OpenStudy (anonymous):

find the domain H(y)=-√ 625-y^4

OpenStudy (anonymous):

Thing about domains, inside radicals, you may not have a negative number :)

OpenStudy (anonymous):

So, we must make it so that \[\huge 625 - y^4 \ge 0\]

OpenStudy (anonymous):

and what is y?

OpenStudy (anonymous):

Solve for it > I suggest factoring :)

OpenStudy (anonymous):

It looks like Peter here gave you the goods to solve it...

OpenStudy (anonymous):

Just remember \[\large a^2 - b^2 = (a+b)(a-b)\] Even more bluntly... \[\large a^4-b^4 = (a^2+b^2)(a+b)(a-b)\]